Equations are given whose graphs enclose a region. Find the area of the region. (Give an exact answer. Do not round.)

f(x) = x2 + 5;  g(x) = −x2; x = 0; x = 1
Answer #1
X = F(x) = X^2 +5 G(x) = -X^2
0 5 0
1 6 -1

Middile of region:

7 6 5 4 3 -F(x) = X^2 + 5 -G(x) = -X^2 2 1 0 0 -1 N